The Eagle Ridge Contractors Association claims the average price of a home in their subdivision is $125,150 with a standard deviation of $7,350. A sample of 36 homes for sale in this subdivision had an average selling price of $123,550. Is there evidence that the costs of homes for sale in this subdivision are actually lower than claimed?
What is the p-value for a one-sided test of the data provided above?

Answer by stanbon(75887)

The Eagle Ridge Contractors Association claims the average price of a home in their subdivision is $125,150 with a standard deviation of $7,350. A sample of 36 homes for sale in this subdivision had an average selling price of $123,550. Is there evidence that the costs of homes for sale in this subdivision are actually lower than claimed?
-----
x-bar = 123550
z(x-bar) = (123550-125150)/[7,350/sqrt(36)] = -1600*6/7350 = -1.306
What is the p-value for a one-sided test of the data provided above?
---
p-value = p(z < -1.306) = normalcdf(-100,-1.306) = 0.0958
---------
Cheers,
Stan H.
---------